# Harnessing quantum back-action for time-series processing

### v1

@@ Abstract
Quantum measurements affect the state of the observed systems via back-action. While projective measurements extract maximal classical information, they drastically alter the system. In contrast, weak measurements balance information extraction with the degree of disturbance. Considering the prevalent use of projective measurements in quantum computing and communication protocols, the potential benefits of weak measurements in these fields remain largely unexplored. In this work, we demonstrate that incorporating weak measurements into a quantum machine-learning protocol known as quantum reservoir computing provides advantages in both execution time scaling and overall performance. We analyze different measurement settings by varying the measurement strength across two benchmarking tasks. Our results reveal that carefully optimizing both the reservoir Hamiltonian parameters and the measurement strength can significantly improve the quantum reservoir computing algorithm performance. This work provides a comprehensive and practical recipe to promote the implementation of weak measurement-based protocols in quantum reservoir computing. Moreover, our findings motivate further exploration of experimental protocols that leverage the back-action effects of weak measurements.
@@ I Introduction
The degree of disturbance induced by a measurement depends on its nature. Conventional projective measurements cause substantial disruption to the quantum system, forcing it into the eigenstate (or a degenerate subspace) associated with the measured outcome. This strong interaction collapses superposition and fully determines the system’s state. In contrast, weak measurements provide a gentler alternative by using indirect interaction between the system and the measurement apparatus [10.1093/acprof:oso/9780199213900.001.0001]. The strength of the coupling and the duration of this interaction govern both the weakness of the measurement and the amount of information extracted. Consequently, weak measurements provide only partial information about the system, inducing a back-action that depends on the degree of measurement weakness. This approach allows the quantum state to retain some coherence, enabling ongoing observation of the system without fully collapsing it. This makes weak measurements a valuable tool for quantum information processing [PhysRevA.62.012307, SINGH2014141]. Weak measurements have been both demonstrated and widely used across various quantum platforms [pan2020weak, spagnolo2022experimental, murch2013observing]. Superconducting qubits, for example, are typically measured through weak measurements using ancillary microwave pulses [naghiloo2019introduction, weber2016quantum]. By adjusting the strength and duration of these pulses, experimenters can tune the measurement strength from weak to strong, with the latter approaching projective measurements in the limit [murch2013observing, Murch2016, doi:10.1126/science.1226897]. In general, the capacity to implement weak measurements relies on the ability to prepare convenient ancillas [Yasuda2023] or to extract information through projectively measuring only a portion of the system [hu2024overcoming, viqueira2023density].
@@
Figure 1: Quantum reservoir computing via weak measurements. (a) Example of the reservoir’s prediction (dashed lines) over the target series (solid lines) on two specific instances ($\eta=5$) for the forward prediction task (top panel) and memory retrieval task (bottom panel) (see section II.2). (b) Schematic view of the information processing step of the QRC algorithm, the coupling strength $g$ between the measurement apparatus and the reservoir is the main parameter of this work’s study. (c) Illustrative example of the effect of weak measurement on a single spin state for different values of the measurement strength, $g$. As $g\to\infty$, the measurement becomes equivalent to a projective measurement, fully collapsing the state onto one of the measurement axes. However, as $g$ decreases, the post-measurement state is no longer collapsed along the measurement axis and is less perturbed. In this analysis, we assume measurements are performed along the $z$-axis.
@@
In this work, we demonstrate that adjusting the strength of quantum measurements enables the optimization of quantum back-action introduced into the system’s evolution, thereby enhancing the performance of machine learning tasks beyond what is achievable without this control. We illustrate our findings within the context of time-series processing using quantum reservoir computing (QRC) [mujal2021opportunities]. QRC belongs to the broader category of quantum machine learning, a field that emerged from the desire to combine machine learning techniques with the unique advantages of quantum systems [dawid2023]. First introduced in [fujii2017harnessing], it is a supervised quantum machine learning technique that leverages the advantages of quantum systems as the exponential growth of their state space with increasing system size. At the same time, it incorporates the simple and efficient training strategy characteristic of classical reservoir computing [jaeger2004harnessing], where the internal dynamics of the reservoir are fixed and only the output layer is trained. Recently, there has been growing interest in QRC and related approaches, with progress both in theoretical advancements [Nakajima2019boosting, Ghosh2019, Ghosh2019prl, Kutvonen2020, martinez2020information, Ghosh2020reconstructing, Tran2020higherorder, Nokkala2021, martinez2021dynamical, Ghosh2021neuromorphic, mujal2021analytical, govia2022nonlinear, Angelatos2021qstatemeas, Nokkala2022, Llodra2022, Domingo2022, Yelin2022, Nokkala2023, GarciaBeni2023, martinez2023finitedim, Krisnanda2023tomographic, mujal2023time, Innocenti2023, Wudarski2023, GarciaBeni2024squeezing, sannia2024dissipation, Nokkala2024, Jaurigue2024memory, Li2024estimating, Kobayashi2024, Domingo2024, Kobayashi2024eco, palacios2024role, Llodra2024] and experimental implementations on diverse platforms [negoro2018machine, Chen2020, Suzuki2022, spagnolo2022experimental, Pfeffer2022, Kubota2023, Yasuda2023, hu2024overcoming, Paternostro2024, Yelin2024]
@@
In QRC, memory refers to the reservoir’s ability to retain and utilize information from past inputs to inform future predictions. It plays a key role in processing temporal series by capturing the time correlations within the dataset. The amount of memory retained by quantum reservoirs has been shown to depend on their dynamical regime [martinez2021dynamical] and is reduced by the effects of quantum back-action [mujal2023time]. When online processing time series, projective measurements introduce significant decoherence, which negatively impacts the memory capacity of the reservoir. As a result, to truly take advantage of the scaling benefits in online processing, a weak measurement approach is more suitable [mujal2023time]. With weak measurements, there is a tradeoff between the amount of information extracted and the quantum back-action introduced. While this reduces disturbance to the reservoir’s memory, it also limits the information obtained, leaving the potential advantage uncertain. This approach has already attracted experimental interest and has been realized in superconducting devices through mid-circuit measurements [hu2024overcoming, Yasuda2023].
@@
Interestingly, in various contexts, controlling the degree of decoherence introduced into the reservoir has been shown to improve its memory capacity [martinez2023finitedim, sannia2024dissipation, monzani2024leveraging, palacios2024role]. In this work, we build on this insight by applying it to a weak measurement-based protocol [mujal2023time]. By leveraging the ability to tune the measurement strength, we optimize the tradeoff between information extraction and back-action. This allows for an online processing scheme that benefits from linear scaling without suffering drawbacks from decoherence. Instead, by carefully controlling the measurement strength, we can enhance the reservoir’s capacity. To achieve this, we introduce a new figure of merit to compare the performance of weak measurement-based protocols, where back-action is controlled, with projective measurement-based protocols that rely on restarting strategies to avoid back-action. By optimizing this metric, we identify the optimal measurement strength that maximizes performance. We illustrate this improvement across different realizations of reservoirs operating in distinct dynamical regimes, validated through two benchmark tasks: the forecasting of the Santa Fe chaotic series [PhysRevA.40.6354] and the short-term memory task . Furthermore, we extend our analysis to a realistic finite resource scenario and show that the enhancement effect is amplified by the favourable scaling properties of the weak measurement protocol.
@@ II Preliminaries
II.1 Quantum reservoir computing via weak measurements
Reservoir computing was proposed as a training-efficient alternative to more complicated recurrent neural network models like Long Short-Term Memory for processing temporal data and time series [jaeger2004harnessing]. Quantum reservoir computing extends the concept by leveraging quantum properties and dynamics, using a quantum system as the reservoir [fujii2017harnessing]. The general pipeline of QRC algorithms is structured in two steps: first the reservoir has to process the time series of interest and then a linear regression model is trained to perform predictions (Fig. 1 (a)). While the latter is a straightforward process, the former includes three stages to be repeated for each element of the dataset: input encoding, time evolution and information extraction (Fig. 1 (b)). At the encoding stage, the information of the $k$-th element of the time series is injected into the amplitudes of the state of a fixed spin of the reservoir [fujii2017harnessing, martinez2021dynamical, mujal2021analytical]. The information is then spread through the system by evolving it for a time $\Delta t$ under a transverse field Ising Hamiltonian $\hat{H}$. Ultimately, we measure a set of observables $\{\langle\hat{O}_{i}\rangle\}$ to extract information from the quantum system. However, when performing projective measurements, it is convenient to reset the system and reprocess the previous $k$ elements without applying measurements before processing the next input, $k+1$. This is because projective measurements erase the stored information in quantum coherences through a destructive process. This procedure, known as the restarting protocol (RSP), is the most common practice in the research community. Recently, a novel protocol was introduced that employs weak measurements instead of projective ones [mujal2023time]. By using weak measurements, only a partial amount of information is lost due to the back-action effect (Fig. 1 (c)), allowing the QRC algorithm to run continuously without restarting at each step in the time series processing. This approach is referred to as the online protocol (OLP) [mujal2023time]. The OLP introduces back-action during the evolution as a tradeoff, allowing for better algorithm scaling by avoiding the need to restart at each input.
@@ II.2 Two benchmark studies
We perform an in-depth analysis of how the back-action effect due to using a weak measurement-based protocol affects the performance of QRC. To do so, we take into account two tasks: the forecasting of a chaotic time series and the efficiency in storing past information. The chaotic time series we use in this work is the Santa Fe laser time series, a common benchmark dataset that follows the chaotic behavior of a $NH_{3}$ laser [PhysRevA.40.6354]. On the other hand, to test the ability to retrieve past information through the short-term memory task, we generate a set of 1000 random points from the uniform distribution in the range $[0,1]$. We build 20 different sub-tasks for both benchmarks where $\eta$, the prediction distance, varies from $1$ to $20$. Given an $\eta$, the target series $k$-th element becomes $y_{k}(\eta)=s_{k\pm\eta}$, where $\vec{s}$ is the original time series. The capacity of the reservoir for each sub-task is evaluated by computing the coefficient of determination $C(\eta)$ of the algorithm predicted series $\vec{\mathtt{y}}$ and the sub-task corresponding true one $\vec{y}(\eta)$:
$C(\eta)=\frac{\text{cov}^{2}(\vec{\mathtt{y}},\vec{y}(\eta))}{\text{var}(\vec{\mathtt{y}})\text{var}(\vec{y}(\eta))}.$ (1)
@@
$C_{\Sigma}=\sum_{\eta=1}^{20}C(\eta).$ (2)
@@ III.1 Reservoir Hamiltonian
In the numerical simulations, we considered systems of $N=6$ spins with random couplings $J_{ij}$ uniformly sorted from the range $[-J/2,J/2]$ and we set $J=1$. Then, when referring to the other parameters of the Hamiltonian we always implicitly express them in units of $J$. At the beginning of the protocol, the initial state of the system is set to be $|\Psi\rangle=|0\rangle^{\otimes N}$.
@@ III.2 Details of the algorithm pipeline
Focusing on the first step of the QRC algorithm pipeline, we provide a detailed description of the framework used in this work. To process the time series, firstly is necessary to normalize it and encode each element in the quantum system. Among the $N$ spins of the reservoir, we consider the following bipartition: we fix one spin to act as the input node, partition $A$, while the others $N-1$ form the partition $B$. For each element of the sequence $s_{k}$, the encoding routine consists of resetting $\hat{\rho}^{A}$ to the pure state $|\psi_{k}\rangle=\sqrt{1-s_{k}}|0\rangle+\sqrt{s_{k}}|1\rangle$ [mujal2021analytical]. After the input injection, the reservoir evolves under $\hat{H}$ for a time $\Delta t=10/J$, $\hat{U}=e^{-i\hat{H}\Delta t}$. In summary, if before the encoding of the element $s_{k}$ the state of the reservoir was $\hat{\rho}_{k-1}$, afterward the reservoir undergoes a completely positive trace-preserving (CPTP) map [fujii2017harnessing, martinez2021dynamical]:
@@
At this point, before the next input injection, it is necessary to extract the elaborated information from the system. As previously described, here we use the online protocol proposed in [mujal2023time] based on weak measurements. The effect of the measurement on the $z$-axis direction is modeled through the action of the matrix $\hat{M}=(\mathbf{I}+e^{-\frac{g^{2}}{2}}\hat{\sigma}_{x})^{\otimes N}$, such that the state after a measurement with strength $g$ is given by:
@@
where $\odot$ indicates the element-wise product. Actually, the CPTP map in this last equation is equivalent to a dephasing channel [martinez2023finitedim]. We can then extract the expectation value of the observable $\hat{\sigma}_{i}^{z}$ on the i-th spin as ${\rm Tr}(\hat{\rho}_{k}\hat{\sigma}_{i}^{z})$. To measure single qubit observables in the $x$ and $y$ direction it is necessary to apply proper rotations to ${\mathcal{L}}_{k}(\hat{\rho}_{k-1})$ before applying $\hat{M}$. For two-spin observables the same description applies.
After having processed all the elements of the sequence $\vec{s}$, we obtain a dataset where for each element of the input data vector we have the corresponding expectation value of the single-spin observables $\langle\hat{\sigma}^{x}_{i}\rangle,\langle\hat{\sigma}^{y}_{i}\rangle,\langle\hat{\sigma}^{z}_{i}\rangle$ for each spin of the reservoir and the two-spin observables $\langle\hat{\sigma}^{x}_{i}\hat{\sigma}^{x}_{j}\rangle,\langle\hat{\sigma}^{y}_{i}\hat{\sigma}^{y}_{j}\rangle,\langle\hat{\sigma}^{z}_{i}\hat{\sigma}^{z}_{j}\rangle$ for each pair of spins of the reservoir. Since measuring in the $x$ and $y$ direction implies applying rotations, this means that the algorithm has to go through the entire dataset three times and not just once.

### v2

@@ Abstract
Quantum measurements affect the state of the observed systems via back-action. While projective measurements extract maximal classical information, they drastically alter the system. In contrast, indirect measurements balance information extraction with the degree of disturbance. Considering the prevalent use of projective measurements in quantum computing and communication protocols, the potential benefits of indirect measurements in these fields remain largely unexplored. In this work, we demonstrate that incorporating indirect measurements into a quantum machine-learning protocol known as quantum reservoir computing provides advantages in both execution time scaling and overall performance. We analyze different measurement settings by varying the measurement strength across two benchmarking tasks. Our results reveal that carefully optimizing both the reservoir Hamiltonian parameters and the measurement strength can significantly improve the quantum reservoir computing algorithm performance. Furthermore, our approach demonstrates improved memory performance when compared with state-of-the-art classical feedback protocols. This work provides a comprehensive and practical recipe to promote the implementation of indirect measurement-based protocols in quantum reservoir computing. Moreover, our findings motivate further exploration of experimental protocols that leverage the back-action effects of indirect measurements.
@@ I Introduction
The degree of disturbance induced by a measurement depends on its nature. Conventional projective measurements cause substantial disruption to the quantum system, forcing it into the eigenstate (or a degenerate subspace) associated with the measured outcome. This strong interaction collapses superposition and fully determines the system’s state. In contrast, indirect measurements provide a gentler alternative by using indirect interaction between the system and the measurement apparatus [10.1093/acprof:oso/9780199213900.001.0001]. The strength of the coupling and the duration of this interaction govern both the disturbance introduced by the measurement and the amount of information extracted. Consequently, indirect measurements can provide only partial information about the system, inducing a back-action that depends on the degree of measurement strength. This approach allows the quantum state to retain some coherence, enabling ongoing observation of the system without fully collapsing it. This makes indirect measurements a valuable tool for quantum information processing [PhysRevA.62.012307, SINGH2014141]. Indirect measurements have been both demonstrated and widely used across various quantum platforms [pan2020weak, spagnolo2022experimental, murch2013observing]. Superconducting qubits, for example, are typically measured through indirect measurements using ancillary microwave pulses [naghiloo2019introduction, weber2016quantum]. By adjusting the strength and duration of these pulses, experimenters can tune the measurement strength from weak to strong, with the latter approaching projective measurements in the limit [murch2013observing, Murch2016, doi:10.1126/science.1226897]. In general, the capacity to implement indirect measurements relies on the ability to prepare convenient ancillas [Yasuda2023] or to extract information through projectively measuring only a portion of the system [hu2024overcoming, viqueira2023density].
@@
Figure 1: Quantum reservoir computing via indirect measurements. (a) Example of the reservoir’s prediction (dashed lines) over the target series (solid lines) on two specific instances ($\eta=5$) for the forward prediction task (top panel) and memory retrieval task (bottom panel) (see section II.2). (b) Schematic view of the information processing step of the QRC algorithm, the coupling strength $g$ between the measurement apparatus and the reservoir is the main parameter of this work’s study. (c) Illustrative example of the effect of indirect measurement on a single spin state for different values of the measurement strength, $g$. As $g\to\infty$, the measurement becomes equivalent to a projective measurement, fully collapsing the state onto one of the measurement axes. However, as $g$ decreases, the post-measurement state is no longer collapsed along the measurement axis and is less perturbed. In this analysis, we assume measurements are performed along the $z$-axis.
@@
In this work, we demonstrate that adjusting the strength of quantum measurements enables the optimization of quantum back-action introduced into the system’s evolution, thereby enhancing the performance of machine learning tasks beyond what is achievable without this control. We illustrate our findings within the context of time-series processing using quantum reservoir computing (QRC) [mujal2021opportunities]. QRC belongs to the broader category of quantum machine learning, a field that emerged from the desire to combine machine learning techniques with the unique advantages of quantum systems [dawid2023]. First introduced in [fujii2017harnessing], it is a supervised quantum machine learning technique that leverages the advantages of quantum systems as the exponential growth of their state space with increasing system size. At the same time, it incorporates the simple and efficient training strategy characteristic of classical reservoir computing [jaeger2004harnessing], where the internal dynamics of the reservoir are fixed and only the output layer is trained. Recently, there has been growing interest in QRC and related approaches, with progress both in theoretical advancements [Nakajima2019boosting, Ghosh2019, Ghosh2019prl, Kutvonen2020, martinez2020information, Ghosh2020reconstructing, Tran2020higherorder, Nokkala2021, martinez2021dynamical, Ghosh2021neuromorphic, mujal2021analytical, govia2022nonlinear, Angelatos2021qstatemeas, Nokkala2022, Llodra2022, Domingo2022, Yelin2022, Nokkala2023, GarciaBeni2023, martinez2023finitedim, Krisnanda2023tomographic, mujal2023time, Innocenti2023, Wudarski2023, GarciaBeni2024squeezing, sannia2024dissipation, Nokkala2024, Jaurigue2024memory, Li2024estimating, Kobayashi2024, Domingo2024, Kobayashi2024eco, palacios2024role, Llodra2024] and experimental implementations on diverse platforms [negoro2018machine, Chen2020, Suzuki2022, spagnolo2022experimental, Pfeffer2022, Kubota2023, Yasuda2023, hu2024overcoming, Paternostro2024, Yelin2024].
@@
In QRC, memory refers to the reservoir’s ability to retain and utilize information from past inputs to inform future predictions. It plays a key role in processing temporal series by capturing the time correlations within the dataset. The amount of memory retained by quantum reservoirs has been shown to depend on their dynamical regime [martinez2021dynamical] and to be reduced by the effects of quantum back-action [mujal2023time]. When online processing time series, projective measurements introduce significant decoherence, which negatively impacts the memory capacity of the reservoir. As a result, to truly take advantage of the scaling benefits in online processing, a weak approach that minimizes the system disturbance is more suitable [mujal2023time]. In this case, there is a tradeoff between the amount of information extracted and the quantum back-action introduced. Although reducing the disturbance preserves the reservoir’s memory, it also limits the information available, leaving the potential advantage uncertain. This approach has already attracted experimental interest and has been realized in superconducting devices through mid-circuit measurements [hu2024overcoming, Yasuda2023].
@@
Interestingly, in various contexts, controlling the degree of decoherence introduced into the reservoir has been shown to improve its memory capacity [martinez2023finitedim, sannia2024dissipation, monzani2024leveraging, palacios2024role]. In this work, we build on this insight by systematically exploring the entire range of indirect measurement regimes and jointly optimizing the measurement strength with the reservoir Hamiltonian parameters, in order to clarify its fundamental role in the reservoir’s performance. By leveraging the ability to tune the measurement strength, we optimize the tradeoff between information extraction and back-action. This allows for an online processing scheme that benefits from linear scaling without suffering drawbacks from decoherence. Instead, by carefully controlling the measurement strength, we can enhance the reservoir’s capacity. To achieve this, we introduce a new figure of merit to compare the performance of the indirect measurement-based protocol, where back-action is controlled, with projective measurement-based protocols that rely on restarting strategies to avoid back-action. By optimizing this metric, we identify the optimal measurement strength that maximizes performance. We illustrate this improvement across different realizations of reservoirs operating in distinct dynamical regimes, validated through two benchmark tasks: the forecasting of the Santa Fe chaotic series [PhysRevA.40.6354] and the short-term memory task . Furthermore, we extend our analysis to a realistic finite resource scenario and show that the enhancement effect is amplified by the favourable scaling properties of the indirect measurement protocol. Finally, we compare our approach with a classical feedback strategy, demonstrating the advantage of exploiting a pure quantum effect such as back-action with respect to incoherent feedback.
@@ II Preliminaries
II.1 Quantum reservoir computing via indirect measurements
Reservoir computing was proposed as a training-efficient alternative to more complicated recurrent neural network models like Long Short-Term Memory for processing temporal data and time series [jaeger2004harnessing]. Quantum reservoir computing extends the concept by leveraging quantum properties and dynamics, using a quantum system as the reservoir [fujii2017harnessing]. The general pipeline of QRC algorithms is structured in two steps: first the reservoir has to process the time series of interest and then a linear regression model is trained to perform predictions (Fig. 1 (a)). While the latter is a straightforward process, the former includes three stages to be repeated for each element of the dataset: input encoding, time evolution and information extraction (Fig. 1 (b)). At the encoding stage, the information of the $k$-th element of the time series is injected into the amplitudes of the state of a fixed spin of the reservoir [fujii2017harnessing, martinez2021dynamical, mujal2021analytical]. The information is then spread through the system by evolving it for a time $\Delta t$ under a transverse field Ising Hamiltonian $\hat{H}$. Ultimately, we measure a set of observables $\{\langle\hat{O}_{i}\rangle\}$ to extract information from the quantum system. However, when performing projective measurements, it is convenient to reset the system and reprocess the previous $k$ elements without applying measurements before processing the next input, $k+1$. This is because projective measurements erase the stored information in quantum coherences through a destructive process. This procedure, known as the restarting protocol (RSP), has long served as the canonical benchmark in the field, since it introduces no additional disturbance to the system. More recently, several alternative protocols have been proposed that focus on offering more practical implementations while still providing good performances [settino, hu2024overcoming, Kobayashi2024, Chen2020], such as the approach based on indirect measurements introduced in Ref. [mujal2023time]. In this scheme, referred to as the online protocol (OLP), only a small amount of information is lost due to the back-action effect (Fig. 1 (c)), allowing the QRC algorithm to run continuously without restarting at each step in the time series processing. The OLP introduces back-action during the evolution as a tradeoff, allowing for better algorithm scaling by avoiding the need to restart at each input.
@@ II.2 Two benchmark studies
We perform an in-depth analysis of how the back-action effect due to using an indirect measurement-based protocol affects the performance of QRC. To do so, we take into account two tasks: the forecasting of a chaotic time series and the efficiency in storing past information. The chaotic time series we use in this work is a subset of $K=2000$ elements of the Santa Fe laser time series, a common benchmark dataset that follows the chaotic behavior of a $NH_{3}$ laser [PhysRevA.40.6354], the exact data points can be found in Ref. [git_tuto]. On the other hand, to test the ability to retrieve past information through the short-term memory task, we generate a set of $K=$ 1000 random points from the uniform distribution in the range $[0,1]$. We build $\eta_{\text{max}}$ different sub-tasks for both benchmarks where $\eta$, the prediction distance, varies from $1$ to $\eta_{\text{max}}$. Given an $\eta$, the target series $k$-th element becomes $t_{k}(\eta)=s_{k\pm\eta}$, where $\vec{s}$ is the original time series. The capacity of the reservoir for each sub-task is evaluated by computing the coefficient of determination $C(\eta)$ of the algorithm predicted series $\vec{p}$ and the sub-task corresponding true one $\vec{t}(\eta)$:
$C(\eta)=\frac{\text{cov}^{2}(\vec{p},\vec{t}(\eta))}{\text{var}(\vec{p})\text{var}(\vec{t}(\eta))}.$ (1)
@@
$C_{\Sigma}=\sum_{\eta=1}^{\eta_{\text{max}}}C(\eta),$ (2)
where $\eta_{\text{max}}$ denotes the $\eta$ for which $C(\eta)\approx 0$.
@@ III.1 Reservoir Hamiltonian
In the numerical simulations, we considered systems of $N=6$ spins with random couplings $J_{ij}$ uniformly sampled from the range [J/2,J/2][-J/2,J/2] and we set $J=1$. Then, when referring to the other parameters of the Hamiltonian we always implicitly express them in units of $J$. At the beginning of the protocol, the initial state of the system is set to be $|\Psi\rangle=|0\rangle^{\otimes N}$.
@@ III.2 Details of the algorithm pipeline
Focusing on the first step of the QRC algorithm pipeline, we provide a detailed description of the framework used in this work. To process the time series, firstly is necessary to normalize it and encode each element in the quantum system. Among the $N$ spins of the reservoir, we consider the following bipartition: we fix one spin to act as the input node, partition $A$, while the others N1N-1 form the partition $B$. For each element of the sequence $s_{k}$, the encoding routine consists of resetting $\hat{\rho}^{A}$ to the pure state $|\psi_{k}\rangle=\sqrt{1-s_{k}}|0\rangle+\sqrt{s_{k}}|1\rangle$ [mujal2021analytical]. After the input injection, the reservoir evolves under $\hat{H}$ for a time $\Delta t=10/J$, $\hat{U}=e^{-i\hat{H}\Delta t}$. The duration of $\Delta t$ is chosen following Ref. [martinez2020information], ensuring that the reservoir exhibits a sufficiently rich dynamical response capable of supporting both linear and nonlinear memory. In summary, if before the encoding of the element $s_{k}$ the state of the reservoir was $\hat{\rho}_{k-1}$, afterward the reservoir undergoes a completely positive trace-preserving (CPTP) map [fujii2017harnessing, martinez2021dynamical]:
@@
At this point, before the next input injection, it is necessary to extract the elaborated information from the system. As previously described, here we use the online protocol proposed in [mujal2023time] based on indirect measurements. These are realized by measuring the momentum quadrature of an ancillary coherent-state pulse coupled to the indirectly measured qubit. As a result, the following measurement operator, e.g. along the $z$-direction, is implemented:
$\hat{\Omega}^{z}_{V}=\frac{1}{\sqrt[4]{2\pi}}\left(e^{-\frac{(V-g)^{2}}{4}}|0\rangle\langle 0|+e^{-\frac{(V+g)^{2}}{4}}|1\rangle\langle 1|\right),$ (5)
where the continuous-valued measurement outcome $V$ and the measurement strength $g$ are in units of the Gaussian width (see Supplementary Information of [mujal2023time] for details). The effect of this measurement on the $z$-axis direction on the unconditional state of the reservoir can be modeled through the action of the matrix $\hat{M}=(\mathbf{I}+e^{-\frac{g^{2}}{2}}\hat{\sigma}^{x})^{\otimes N}$, such that the state after a measurement with strength $g$ is given by:
@@
where $\odot$ indicates the element-wise product. Actually, the CPTP map in this last equation is equivalent to a dephasing channel [martinez2023finitedim]. We can then extract the expectation value of the observable $\hat{\sigma}_{i}^{z}$ on the i-th spin as ${\rm Tr}(\hat{\rho}_{k}\hat{\sigma}_{i}^{z})$. To measure single qubit observables in the $x$ and $y$ direction it is necessary to apply proper rotations to ${\mathcal{L}}_{k}(\hat{\rho}_{k-1})$ before applying $\hat{M}$. For two-spin observables the same description applies. Since the RSP is equivalent to dynamics without back-action, it can be conveniently simulated by setting $g=0$ in this model, for which $M$ reduces to an all-ones matrix and has no effect on the state.
After having processed all the elements of the sequence $\vec{s}$, we obtain a dataset where for each element of the input data vector we have the corresponding expectation values of the single-spin observables $\langle\hat{\sigma}^{x}_{i}\rangle,\langle\hat{\sigma}^{y}_{i}\rangle,\langle\hat{\sigma}^{z}_{i}\rangle$ for each spin in the reservoir, as well as the two-spin observables $\langle\hat{\sigma}^{x}_{i}\hat{\sigma}^{x}_{j}\rangle,\langle\hat{\sigma}^{y}_{i}\hat{\sigma}^{y}_{j}\rangle,\langle\hat{\sigma}^{z}_{i}\hat{\sigma}^{z}_{j}\rangle$ for all pairs of spins $i\neq j$. Single-spin variances, i.e., $\langle(\hat{\sigma}^{\alpha}_{i})^{2}\rangle,$ are also computed as a consistency check, though they do not affect the performance. Since measuring in the $x$ and $y$ direction implies applying rotations, this means that the algorithm has to go through the entire dataset three times and not just once.
From the obtained $(K\times 81)$ dataset, we discard the first $20$ rows, which correspond to the expectation values of the initial $20$ elements of the time-series. This washout step and the fading-memory property of the reservoir ensure that its state becomes independent of the initial conditions and depends only on the recent input history. The choice of $20$ steps follows the analysis of the required washout length for this system reported in Ref. [martinez2020information].